# Large Language Models powered Network Attack Detection: Architecture, Opportunities and Case Study

### v1

@@ Abstract
Network attack detection is a pivotal technology to identify network anomaly and classify malicious traffic. Large Language Models (LLMs) are trained on a vast corpus of text, have amassed remarkable capabilities of context-understanding and commonsense knowledge. This has opened up a new door for network threat detection. Researchers have already initiated discussions regarding the application of LLMs on specific cyber-security tasks. Unfortunately, there is still a lack of comprehensive elaboration how to mine LLMs’ potentials in network threat detections, as well as the opportunities and challenges. In this paper, we mainly focus on the classification of malicious traffic from the perspective of LLMs’ capability. We present a holistic view of the architecture of LLM-powered network attack detection, including Pre-training, Fine-tuning, and Detection. Especially, by exploring the knowledge and capabilities of LLM, we identify three distinct roles LLM can act in network attack detection: Classifier, Encoder, and Predictor. For each of them, the modeling paradigm, opportunities and challenges are elaborated. Finally, we present our design on LLM-powered DDoS detection as a case study. The proposed framework attains accurate detection on carpet bombing DDoS by exploiting LLMs’ capabilities in contextual mining. The evaluation shows its efficacy, exhibiting a nearly $35$% improvement compared to existing systems.
@@
network attack detection, malicious traffic classification, Large Language Model, .
@@ I Introduction
Network attack detection is the cornerstone of network security since it is about the classification and identification network intrusion or malicious traffic. It usually oversees network behavior and detects network attacks that might compromise the security, integrity, and functionality of networks. According Cloudflare reports [1], each day in 2024 has $209B$ network attacks with $53\%$ increased attacks. The largest attack traffic is recorded with exceeding $5.6Tbps$ in one DDoS attack. It remains an urgent issue for academia and industry to detect and prevent network threats timely and precisely.
@@
With the increased volume of network traffic and evolving natures of attacks, the existing attacks classification methods, such as rule-based, machine-learning or deep-learning and anomaly-based methods, face several issues [2]. Rule-based approach has been extensively embraced by the industry for its practicality and effectiveness. But the hand-crafted rules are time-consuming and require a great deal of expertise. Supervised machine-learning or deep-learning methods require a amount of labeled data for the training, which is impractical due to the evolving network traffic and labor-intensive labeling work. Unsupervised anomaly detection methods can detect unknown attacks by identifying the pattern that deviate from normal network traffic, but it faces the issues of pattern drift.
@@
Recently, Large Language Models (LLMs) [3] have demonstrated powerful potentials in natural language processing and computer vision, such as language understanding, image classification, and multi-modal learning, etc. LLMs, including BERT, GPT series, etc. are based on the Transformer neural network architectures [4, 5], and learn the in-context pattern of language from a vast corpus of text using self/semi-supervised learning techniques, The key capabilities of LLMs on context-understanding and embedded vast knowledge open a new door for network attack detections [6].
@@
Unfortunately, there is still a lack of comprehensive elaboration on the current LLM-powered network attack detections, its overall architecture, as well as the opportunities and challenges therein. A few of surveys [7, 8] have summarize the recent researches on the application of LLMs in offensive and defensive cyber-security as well as cyber-ethics. Most of them scatter the discussions on various cyber-security topics, not delving into the technical architecture and illustrating a clear picture of how LLMs identify malicious network traffic.
@@
In this paper, we fill this gap from the perspective of LLMs’ capability. We present an architecture of LLM-powered network attack detection. Especially, by exploring the potential of LLM, we introduce three roles of LLMs can act in network attack detection, Classifier, Encoder, and Predictor. For each role, the detection paradigm, opportunities and challenges are elaborated. We try to answer the three question. 1) How to exploit LLMs’ potentials in network attack detections. 2) How to enable LLMs to understand non-language network traffic. 3) What capability LLMs can provide in the detection. To the best of our knowledge, this paper is one of the first works that systematically present the architecture and paradigm of LLM-powered network attack detection.
@@
  • We present an holistic view of the architecture of LLM-powered network attack detection, from the perspective of exploring LLMs’ potentials.
  • We point out the roles LLM can play in network attack detection, and corresponding opportunities and challenges.
@@
The rest of the article is organized as follows: Sec. II present a holistic view of the architecture of LLM-powered network attack detection. In Sec. III, the LLM’s roles, opportunities and challenge are introduced in details. Sec. IV demonstrates a case study of LLM-power DDoS detection. Finally, the conclusion and future work are given in Sec. V.
II The Architecture of LLM-powered Network Attack Detection
@@
In this section, we introduce the architecture of LLM-powered network attack detection. As in Fig. 1, it consists of three stages, i.e., Pre-training, Task Training, and Detection.
Figure 1: Architecture of LLM-powered Network Attack Detection.
II-A Unsupervised Pre-training
@@
LLM pre-training is a crucial technology of training a model with transfer learning capability. It is a self-supervised learning method, which means the model learns from the data itself without relying on explicit human-labeled supervision. The aim is to enable the model to learn the general rules and semantic knowledge of the sequential data, such as the relationships between words, the structure and semantics of sentences, etc. The pre-trained models can be fine-tuned in various downstream tasks with improved performance and reducing the demand for a large amount of labeled data.
There are three parts of LLM pre-training for network attack detection, Data Preparation, Masked Traning and Auto-regressive Training.
@@
Data Preparation: Pre-training data for an LLM in network attack detection could include text corpus and network data. A large amount of text data, such as academic papers, books, web pages, etc. allow the model to learn general language patterns and vast knowledge. Network data include packet-level, flow-level and topology-level which make LLM learn network communication pattern.
@@
The pre-trained model, also called foundational model, can be trained with network specific data. But recently researches find [3] that the model trained with nature-language text is able to understand multi-model data. Thus, LLM foundation model pre-trained with text data will be a proper choice for network attack detection [7].
@@
Masked Training: Masked pre-training is trained to predict the masked token in a sequence. The representative model is developed by Google, Bidirectional Encoder Representations from Transformers(BERT) [4]. The model is presented with a large amount of data where certain tokens are masked. The LLM’s task is to predict these masked tokens based on the surrounding context. This training helps the LLM to understand context within data.
@@
Auto-regressive Training: The model is trained to predict the next token in a sequence given the previous tokens. The well-known model of this type is Generative Pre-trained Transformer(GPT) [5], a series of pre-trained language models developed by OpenAI. When applied to network attacks detection, the LLM can learn the sequential patterns of traffic.
II-B Task-specific Training
@@
Task-specific training is a crucial step in adapting a pre-trained foundation model to a specific task, such as traffic classification, anomaly detection, etc. It refines the general knowledge acquired during pre-training, enabling the LLM to better handle the specific tasks in network attack detections. There are two main types of task training, In-Context Learning and Fine-tuning.
@@
In-Context Learning: In-Context Learning (ICL) allows the LLM to learn and perform tasks based on examples provided within the input context, Prompt, without explicitly updating the model’s parameters. In the context of network attack detection, a prompt could be a description of a network security scenario, along with some examples of benign and malicious activities.
@@
ICL is powerful as it enables the model to quickly adapt to new tasks without the need for extensive training, making it efficient for handling time-changing network attacks.
@@
Task Fine-tuning: Task Fine-tuning focuses on training the LLM for a specific task by supervised learning with labeled training data, as shown in Fig. 1. The pre-trained model are used as a starting point whose parameters are updated by supervised gradient. This significantly reduces the amount of data and time required for the new task training from scratch.
@@
The strength of this approach is targeted to optimize the model’s performance with high accuracy. But the drawback is that a lot of labeled data for training is needed, and updating model parameter is computation-consuming.
II-C Detection/Inferencing
@@
According to the potential capability of LLM, it can perform three crucial roles in network attack detection: Classifier, Encoder, and Predictor. Each role contributes uniquely to enhancing the effectiveness and efficiency.
Classifier: Exploiting the commonsense knowledge of LLM, it is possible for LLM to serve as an Intelligent Agent in network attack detection [5]. It is tasked with categorizing network traffic into different classes, such as benign traffic, malicious traffic, or specific types of threats.
@@
As in Fig. 1, The LLM is fed into Prompt, a text description about task introduction, some examples of network flow description, and your questions. The LLMs answer whether the flow is benign or malicious. It applies the knowledge it has learned during pre-training, ICL or fine-tuning to make a classification decision.
@@
Encoder: Exploiting the capability of context understanding, LLM often act as a feature extractor, i.e., representation learner [9, 10]. As an encoder, the LLM extracts meaningful representations or features from network traffic, often in the form of a vector. This encoded representation captures the essential features and relationships within the data.
@@
Comparing to the traditional method of human feature engineering, the encoding process by the LLM helps in condensing complex network data into a format that can be easily processed, facilitating more efficient attack detection.
@@
Predictor: By exploiting the capability of LLMs on masked-token prediction or next-token auto-regressively generation, LLM can predict future traffic pattern [11]. As a predictor, the LLM aims to forecast future activities or trends related to network traffic. It uses historical data, current network status information, and the knowledge it has acquired during training to make predictions.
@@ III Opportunities and Challenges
In this section, we discuss the roles, research opportunities and challenges in LLM-powered network attack detection. Specifically, we focus on the detailed diagram of LLM as Classifier, Encoder, and Predictor.
@@ III-A LLM as Classifier
Inspired by the concept of LLM agents, where an agent is trained to handle planning, decision-making, and control, the idea of making an LLM an intelligent agent for network attack detection is promising [12]. The network traffic data and carefully-crafted Prompts are feed into the LLM. It classifies network traffic by the guidance provided in the Prompts, leveraging its pre-trained language-understanding knowledge and pattern-recognition capabilities. The wealth human-knowledge in LLM provides great opportunities for network attack detection, but also comes with challenges.
As showed in Fig. 2, the main components of LLM-based classifier consists of three crucial modules: Data Processing, Prompt Construction, and fine-tuned LLM. The modules work together to make LLM understand network traffic and inference which flow is malicious.
Data Processing. Before feeding into the LLM, raw network traffic first needs to be processed to a form that the LLM can understand. There are two main processing methods: text description and embedding learning.
Text Description: Text description involves converting raw network traffic data into language text. For example, network packet headers, which contain details like source and destination IP addresses, port numbers, and protocol types, can be translated into sentences. This text-based representation allows the LLM to utilize its natural-language-processing capabilities to analyze the network data. But it has limitations in handling numerical data with high precision, since text could not capture complex patterns in numerical data.
Embedding Learning: Embedding learning aims to transform the raw network traffic data into a numerical representation, often in the form of vectors or tokens. For network packet data, algorithms can be used to extract features such as the frequency of different protocol types, the distribution of packet sizes, and the relationship between source and destination IP addresses. These features are then mapped into a vector space. This numerical representation can be more easily processed by the LLM, enabling it to perform operations such as similarity comparison and pattern recognition more efficiently.
@@
Here is an example of Prompt. It consists of three parts:
Task Description: This part clearly defines the objective and instruction of the network attack detection task. By providing a clear task description, the LLM is better able to focus its analysis on the relevant aspects of the network traffic.
Few-shot Samples: Usually there are a small number of labeled examples to illustrate the input data and their corresponding attack classifications. For instance, a few-shot sample could include an example of benign network traffic with a description of why it is considered benign. These samples help the LLM learn the patterns and features associated with the Malicious/Benign labels. They are equivalent to the training data set in supervised learning.
Format Statement: It specifies the expected output format of the LLM detection. For example, ”Please respond with either ’Benign’ or ’Malicious’ to indicate the nature of the network flows, and provide a brief explanation for your classification.” This ensures that the LLM’s response is in a consistent format without nonsensical generated text.
Fine-tuned LLM. General LLM should be fine-tuned to know how to detect network attack as mentioned in II. Two types of task-specific model adaptation are both applicable to network attack detection: Fine-tuning and In-Context Learning.
Fine-tuning is a supervised learning which needs a few of labeled samples. As the results validated in  [13], it can achieve higher precisions. But it requires modifying the LLM’s parameters through supervised gradient. This will allow the LLM to better adapt to the specific requirements of network attack detection tasks.
@@
Although LLM classifiers exhibit amazing potentials in network attack detection, there are still many non-intuitive key challenges that need to be overcome.
@@
Hallucination: LLMs are probability models of language tokens, prone to hallucination anyway. This may generate false positive results in the context of network attack detection. For example, the LLM might misinterpret benign network traffic fluctuations as signs of a malicious attack. This could be due to its over-reliance on certain patterns in the training data.
Computation Consumption: Network attack detection often demands real-time or near-real-time analysis of traffic. However, as the volume of network traffic increases, the inferencing time taken by the LLM may exceed the acceptable latency for effective threat detection.

### v2

# Large Language Models powered Malicious Traffic Detection: Architecture, Opportunities and Case Study

@@ Abstract
Malicious traffic detection is a pivotal technology for network security to identify abnormal network traffic and detect network attacks. Large Language Models (LLMs) are trained on a vast corpus of text, have amassed remarkable capabilities of context-understanding and commonsense knowledge. This has opened up a new door for network attacks detection. Researchers have already initiated discussions regarding the application of LLMs on specific cyber-security tasks. Unfortunately, there remains a lack of comprehensive analysis on harnessing LLMs for traffic detection, as well as the opportunities and challenges. In this paper, we focus on unleashing the full potential of Large Language Models (LLMs) in malicious traffic detection. We present a holistic view of the architecture of LLM-powered malicious traffic detection, including the procedures of Pre-training, Fine-tuning, and Detection. Especially, by exploring the knowledge and capabilities of LLM, we identify three distinct roles LLM can act in traffic classification: Classifier, Encoder, and Predictor. For each of them, the modeling paradigm, opportunities and challenges are elaborated. Finally, we present our design on LLM-powered DDoS detection as a case study. The proposed framework attains accurate detection on carpet bombing DDoS by exploiting LLMs’ capabilities in contextual mining. The evaluation shows its efficacy, exhibiting a nearly $35$% improvement compared to existing systems.
@@
network intrusion detection, malicious traffic detection, traffic classification, Large Language Model, DDos.
@@ I Introduction
Malicious traffic detection is the cornerstone of network security since it is about the classification and identification of network attack traffic. It usually oversees network behavior and detects network anomalies that might compromise the security, integrity, and functionality of networks. According to Cloudflare reports [1], each day in 2024 witnessed 209 billion network attacks events, marking a 53% annual increase. The largest recorded DDoS attack in 2024 even surpassed 5.6 terabits per second in traffic volume. It remains critical challenges for both academia and industry to detect and prevent network attacks timely and precisely.
@@
With the exponential growth of network traffic and evolving natures of attacks, existing malicious traffic classification methods, such as rule-based, machine-learning or deep-learning and anomaly-based methods, are confronting critical challenges [2]. Rule-based approach has been extensively embraced by the industry for its practicality and effectiveness. But the hand-crafted rules are time-consuming and require a great deal of expertise. Supervised machine-learning [3] or deep-learning methods [4] require a amount of labeled data for the training, which is impractical due to the labor-intensive labeling work and huge amounts of network traffic. Unsupervised anomaly detection methods [2] can detect zero-day attacks by identifying the pattern that deviate from normal network traffic, but it faces the issues of traffic drift.
@@
Recently, Large Language Models (LLMs) [5] have demonstrated powerful potentials in natural language processing and computer vision, such as language understanding, image classification, and multi-modal learning, etc. LLMs, including BERT [6], GPT [7] series, etc. are based on the Transformer neural network architectures, and learn the in-context pattern of token sequence from a vast corpus of text using self/semi-supervised learning techniques, The key capabilities of LLMs on context-understanding and embedded vast knowledge open a new door for network attack detections [8].
@@
Unfortunately, there is still a lack of comprehensive elaboration on the current LLM-powered malicious traffic detections, its overall architecture, as well as the opportunities and challenges therein. A few of surveys [9, 10] have summarize the recent researches on the application of LLMs in offensive and defensive cyber-security. Most of them exploit LLMs’ capability on nature language processing such as network intent understanding, network operation, security knowledge Q&A, etc. Few of them explore LLMs in traffic classification. They don’t delve into the technical architecture and illustrating a clear picture of how LLMs enable malicious traffic detection.
@@
In this paper, we fill this gap from the perspective of LLMs’ capability. We present an architecture of LLM-powered malicious traffic detection. Especially, by exploring the potential of LLM, we introduce three roles of LLMs can act in traffic classification, Classifier, Encoder, and Predictor. For each role, the detection paradigm, opportunities and challenges are elaborated. We try to answer the three question. 1) How to exploit LLMs’ potentials in malicious traffic detections. 2) How to enable LLMs to understand non-language network traffic. 3) What capability LLMs can provide in the detection. To the best of our knowledge, this paper is one of the first works that systematically present the architecture and paradigm of LLM-powered malicious traffic detection.
@@
  • We present an holistic view of the architecture of LLM-powered malicious traffic detection, from the perspective of exploring LLMs’ potentials.
  • We point out the roles LLM can play in traffic classification, and corresponding opportunities and challenges.
@@
The rest of the article is organized as follows: Sec. II present a holistic view of the architecture of LLM-powered malicious traffic detection. In Sec. III, the LLM’s roles, opportunities and challenge are introduced in details. Sec. IV demonstrates a case study of LLM-power DDoS detection. Finally, the conclusion and future work are given in Sec. V.
II The Architecture of LLM-powered Malicious Traffic Detection
@@
In this section, we introduce the architecture of LLM-powered malicious traffic detection. As in Fig. 1, the framework comprises three sequential stages, i.e., Pre-training, Task Training, and Detection.
Figure 1: Architecture of LLM-powered Malicious Traffic Detection.
II-A Self-supervised Pre-training
@@
LLMs pre-training is a crucial technology of training a model with transfer learning capability. It is a self-supervised learning method, which means the model learns from the data itself without relying on explicit human-labeled supervision. The aim is to enable the model to learn the general rules and semantic knowledge of the sequential data, such as the relationships between data, the structure and semantics of sequences, etc. The pre-trained models can be fine-tuned in various downstream tasks with improved performance and reducing the demand for a large amount of labeled data.
@@
Pre-training Data for LLM-powered malicious traffic detection could include text corpus and network data. A large amount of text data, such as academic papers, books, web pages, etc. allow the model to learn general language patterns and vast knowledge. Network data include packet-level, flow-level and topology-level network traffic data which make LLM learn network communication pattern.
@@
Pre-trained Models, also called foundational model, can be trained with network specific data. But recent studies [5] have revealed that the model trained on nature-language text is able to understand multi-model data. Thus, LLM foundation model pre-trained on text data will be a proper choice for malicious traffic detection.
Pre-training Methods can be primarily categorized into two paradigms: Masked Training and Auto-regressive Training, each employing distinct learning mechanisms to enable model generalization:
@@
Masked Training: Masked pre-training involves training models to predict masked tokens in sequences. The canonical model is developed by Google, Bidirectional Encoder Representations from Transformers(BERT) [6]. The model is presented with a large amount of data where certain tokens are masked. The LLM’s task is to predict these masked tokens based on the surrounding context. This training helps the LLM to understand context within data.
@@
Auto-regressive Training: The model is trained to predict the next token in a sequence given the previous tokens. The well-known model of this type is Generative Pre-trained Transformer(GPT) [7], a series of pre-trained language models developed by OpenAI. The model processes sequences in a left-to-right manner, using a causal attention mechanism. This enables it to learn the sequential patterns of network traffic in malicious traffic detection.
II-B Task-specific Post-training
@@
Task-specific post-training is a crucial step in adapting a pre-trained foundation model to a specific task, such as traffic classification, anomaly detection, etc. It refines the general knowledge acquired during pre-training, enabling LLM to develop task-specific capabilities for malicious traffic detection. Specifically, task-specific post-training can be categorized into two primary paradigms: In-Context Learning and Fine-tuning.
@@
In-Context Learning: In-Context Learning (ICL) [11] allows the LLM to learn from the examples provided within the input context, Prompt, without explicitly updating the model’s parameters. In the context of malicious traffic detection, this paradigm enables zero-shot/few-shot inference with a prompt consisting of a description of a network security scenario, along with some examples of benign and malicious activities.
@@
ICL is powerful as it enables the model to quickly adapt to new tasks without requiring extensive re-training LLM models, making it highly efficient for addressing time-evolving network attacks. But it falls short of detection accuracy.
@@
Fine-tuning: Task Fine-tuning focuses on training the LLM for a specific task by supervised learning with labeled training data, as shown in Fig. 1. The pre-trained model are used as an initialization whose parameters are updated by supervised gradient. This significantly reduces the amount of data and time required compared to training models from scratch.
@@
The advantage of this approach [9] is to optimize the model’s performance with high accuracy. But the drawback is that a lot of labeled data for training is needed, and updating model parameter is computation-consuming.
II-C Detection
@@
According to the potential capability of LLM, it can perform three crucial roles in malicious traffic detection: Classifier, Encoder, and Predictor. Each role contributes uniquely to enhancing the effectiveness and efficiency.
@@
Classifier: Exploiting the commonsense knowledge of LLM, it is possible for LLM to serve as an intelligent Agent [7] in malicious traffic detection. As in Fig. 1, The LLM is fed into Prompt, a text description about task introduction, some examples of network flow description, and your questions. The LLMs answer whether the flow is benign or malicious. It applies the knowledge it has learned during pre-training and few-shots to make a classification decision.
@@
Encoder: Exploiting the capability of context understanding, LLM often act as a feature extractor [12], i.e., representation learner. As an encoder, the LLM extracts meaningful representations or features from network traffic, often in the form of a vector. This encoded representation captures the essential features and relationships within the network traffic.
@@
Comparing to the traditional method of human feature engineering, the encoding process by the LLM helps in condensing complex traffic data into a format that can be easily processed, facilitating more efficient threat detection.
@@
Predictor: By exploiting the capability of LLMs on masked-token prediction or next-token auto-regressively generation [13], LLM can predict future traffic pattern. As a predictor, the LLM aims to forecast future activities or trends related to network traffic. It uses historical data, current network status information, and the knowledge it has acquired during training to make predictions.
@@ III Opportunities and Challenges
In this section, we discuss the roles, research opportunities and challenges in LLM-powered malicious traffic detection in details. Specifically, we focus on the novel diagram of LLM as Classifier, Encoder, and Predictor.
@@ III-A LLM as Classifier
Inspired by the concept of LLM agent, where an agent is trained to handle planning, decision-making, and control, the idea of making an LLM an intelligent agent for malicious traffic detection is promising [11]. The network traffic data and carefully-crafted Prompts are feed into the LLM. It classifies network traffic by the guidance provided in the Prompts, leveraging its pre-trained language-understanding knowledge and pattern-recognition capabilities. The wealth human-knowledge in LLM provides great opportunities for malicious traffic detection, but also comes with challenges.
As showed in Fig. 2, the main components of LLM-based classifier consists of three crucial modules: Traffic Processing, Prompt Construction, and LLM instruction fine-tuning or ICL. The modules work together to make LLM understand network traffic and inference which flow is malicious.
Traffic Processing. Before feeding into the LLM, raw network traffic first needs to be processed to a form that the LLM can understand. There are two main processing methods: text description and embedding learning.
Text Description: Text description involves converting raw network traffic data into language text. For example, network packet headers, which contain details like source and destination IP addresses, port numbers, and protocol types, can be translated into texts. This text-based representation allows the LLM to utilize its natural-language-processing capabilities to analyze the network data. But it has limitations in handling numerical data with high precision, since text could not capture complex patterns in numerical data.
Embedding Learning: Embedding learning aims to transform the raw network traffic data into a numerical representation, often in the form of vectors or tokens. For network traffic data, algorithms can be used to extract features such as the frequency of different protocol types, the distribution of packet sizes, and the relationship between source and destination IP addresses. These features are then mapped into a vector space. This numerical representation can be more easily processed by the LLM, enabling it to perform operations such as similarity comparison and pattern recognition more efficiently.
@@
Here is an example of Prompt template. It includes:
Task Description: This part clearly defines the objective and instruction of the network threat detection task. By providing a clear task description, the LLM is better able to focus its analysis on the relevant aspects of the network traffic.
Few-shot Samples: Usually there are a small number of labeled examples to illustrate the input data and their corresponding traffic classifications. For instance, a few-shot sample could include an example of benign network traffic with a description of why it is considered benign. These samples help the LLM learn the patterns and features associated with the Malicious/Benign labels. They are equivalent to the training data set in supervised learning.
Output Statement: It specifies the expected output format of the LLM detection. For example, Please respond with either ’Benign’ or ’Malicious’ to indicate the nature of the network flows, and provide a brief explanation for your classification. This ensures that the LLM’s response is in a consistent format without nonsensical generated text.
ICL or Fine-tuning. LLM should be fine-tuned to know how to detect malicious traffic as mentioned in II. Two types of task-specific model adaptation are both applicable to traffic detection: Fine-tuning and In-Context Learning.
Instruction fine-tuning is a supervised learning which needs a few of labeled texture samples. As the results validated in [11], it can achieve higher precisions. But it requires modifying the LLM’s parameters through supervised gradient. This will allow the LLM to better adapt to the specific requirements of network threat detection tasks.
@@
Although LLM classifiers exhibit amazing potentials in malicious traffic detection, there are still many non-intuitive key challenges that need to be overcome.
@@
Hallucination: LLMs are probability models of language tokens, prone to hallucination anyway. This may generate false positive results in the context of malicious traffic detection. For example, the LLM might misinterpret benign network traffic fluctuations as signs of a malicious attack. This could be due to its over-reliance on certain patterns in the training data.
Computation Consumption: Malicious traffic detection often demands real-time or near-real-time analysis of traffic. However, due to the computation-consumption of LLM, it may takes LLM to inference with unacceptable latency.